OpenStudy (anonymous):

I GIVE MEDALS. A motorcycle is capable of accelerating at 5.1 m/s2. Starting from rest, how far can it travel in 1.5 seconds? A. 3.8 m B. 4.5 m C. 3.2 m D. 5.7 m

OpenStudy (anonymous):

D

OpenStudy (anonymous):

Abby I see you post these questions a lot, do you take any time to try to learn them/ think about them? if it can accelerate 5.1 m/s^2 in only one second, how could it be any other answer other than D? the rest are less than 5 even.

OpenStudy (anonymous):

linear displacement of a particle with time is given by s=ut +1/2at^2 u->initial velocity since u=0 s=1/2at^2 s=1/2*5.1*1.5*1.5

OpenStudy (anonymous):

Yeah, actually. i'm studying for finals so im going over which is why there are so many questions. thanks for so being concerned. @robertsn & thanks @MayankD for helping and not taking time to figure out what I learn.

OpenStudy (anonymous):

& is s=ut+1/2at^2 always the equation for linear displacement? @MayankD

OpenStudy (anonymous):

:) here to help

OpenStudy (anonymous):

yes only when the accleration is constant in magnitude and direction